OpenStudy (anonymous):

A 1.000 g sample of a metal chloride, MCl2, is dissolved in water and treated with excess aqueous silver nitrate. The silver chloride that formed weighed 1.286g. Calculate the atomic mass of M.

OpenStudy (anonymous):

given 1 gm of MCl2 fors 1.286g of AgCl molar mass of AgCl is 143.5 g 1.286g of AgCl contains 1.286/143.5 mols of AgCl =0.0089616 mols each mol of MCl2 release two Cl- ions hence number of mols MCl2 given(1gm) is 0.0089616 /2=0.0044808 mols hnce one mol of MCl2 weighs 1/0.0044808= 223.17g